OpenStudy (anonymous):

Solve: 1/m=a/b a. m=ab b. m=a/b c. m=b/a

OpenStudy (asnaseer):

In any equation of the form: something = something_else what ever you do to the left-hand-side of the equation also has to be done to the right-hand-side of the equation. with this in mind, what would you need to do to the left-hand-side of your equation to get from 1/m to m?

OpenStudy (anonymous):

multiply by 1?

OpenStudy (asnaseer):

that would just give:\[1\times\frac{1}{m}=\frac{1}{m}\]

OpenStudy (asnaseer):

have you heard of "inversion"?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

so would you multiply it by a/b too?

OpenStudy (asnaseer):

ok, given some fraction like:\[\frac{a}{b}\]we can invert it to get:\[\frac{b}{a}\]so given the fraction:\[\frac{1}{m}\]if we invert it, what do you think we will get?

OpenStudy (asnaseer):

remember inversion just involves swapping the numerator and denominator of the fraction

OpenStudy (anonymous):

m/1

OpenStudy (asnaseer):

perfect!

OpenStudy (asnaseer):

so now remember the rule I gave you at the start - what ever you do to the left-hand-side of the equation also has to be done to the right-hand-side of the equation. using this rule, we know that since we inverted the left-hand-side, we must also now invert the right-hand-side

OpenStudy (asnaseer):

so what would you get if you inverted:\[\frac{a}{b}\]

OpenStudy (anonymous):

m/1=b/a?

OpenStudy (asnaseer):

that is correct - well done! :)

OpenStudy (asnaseer):

and remember that m/1 is the same as just m.

OpenStudy (anonymous):

okay so the answer would be b/a, correct?

OpenStudy (asnaseer):

yes :)

OpenStudy (anonymous):

thank you!! :)

OpenStudy (asnaseer):

yw :)